OpenStudy (anonymous):

Help with Tan^-1(sin(2pi/3))

OpenStudy (anonymous):

Draw the right triangle.

OpenStudy (anonymous):

Make sure you notice which quadrant 2π/3 is in.

OpenStudy (p0sitr0n):

sin(2pi/3)=2sin(pi/3)cos(pi3)

OpenStudy (p0sitr0n):

*cos(pi/3)

OpenStudy (p0sitr0n):

2*1/2*\[\sqrt{3}/2 \] = 3^1/2/2

OpenStudy (anonymous):

sin(2π/3) = sin(π/3)

OpenStudy (anonymous):

*cough* unit circle *cough*

OpenStudy (anonymous):

I was thinking Sin(2pi/3) would be in the II quadrant.

OpenStudy (anonymous):

I see it on the unit circle, but my teacher told me, Sine can't be in the quadrant because it has a range of +/- pi/2

OpenStudy (p0sitr0n):

range of sine is from -1 to 1. its domain is R

OpenStudy (anonymous):

Oh, so the range of sin^-1 is +/- pi/2 then right?

OpenStudy (p0sitr0n):

its an inverse function. dom f(x) = range f^-1(x)

OpenStudy (p0sitr0n):

range of arcsin (sin^-1) is R. its like the sin but tilted 90 degrees. domain is [-1,1]

OpenStudy (anonymous):

You'll only have to worry about the domain of arctangent here, but it won't be a problem.

OpenStudy (anonymous):

Okay, I got that part down, so how should I figure out the rest of the problem? Should my answer be in radians or...

OpenStudy (anonymous):

\[\sin(\frac{2\pi}{3})=\frac{\sqrt{3}}{2}\] \[\tan^{-1}(\frac{\sqrt{3}}{2})\] is a calculator exercise

OpenStudy (anonymous):

But the thing is, the teacher doesn't want us to use a calculator. It's more of a memorize the unit circle excercise.